Jaydybird:

Two linear functions are represented in different formats. Function 1: x y −2 ​ −1 ​ 0 ​1 2 ​3 5 ​6 Function 2: Graph: (-1,-4) (2,7) Which statements are true? Select each correct answer. Function 1 has a greater rate of change than function 2. Function 2 has a greater rate of change than function 1. Function 1 has a greater y-intercept than function 2. Function 2 has a greater y-intercept than function 1.

jimthompson5910:

Function 1 has the two points (-2,-1) and (0,1) Do you know how to find the slope of the line through these two points?

Jaydybird:

no

TWALKER16:

do you need help?

Jaydybird:

yes please!

TWALKER16:

ok

jimthompson5910:

First you subtract the y coordinates: -1-1 = -2 Then subtract the x coordinates in the same order -2-0 = -2 Lastly you divide the two results (-2)/(-2) = 1 So the slope of the first function is 1


Or you could use the slope formula \(\large m = \frac{y_2 - y_1}{x_2-x_1}\) \(\large m = \frac{1-(-1)}{0-(-2)}\) \(\large m = \frac{1+1}{0+2}\) \(\large m = \frac{2}{2}\) \(\large m = 1\) The slope formula says the same basic idea as the previous section where we subtracted the y values, then the x values, and divided the differences

jimthompson5910:

Hopefully that makes sense. If so, then please tell me what the slope of function 2 would be.

Jaydybird:

so would it be m=-11/-3 ?

jimthompson5910:

that can be simplified further

Jaydybird:

so then it would be 3 1/3 ?

TWALKER16:

Yes

jimthompson5910:

11/3 is not the same as 3 & 1/3

Jaydybird:

ya thats what i meant

Jaydybird:

wait now im confused

jimthompson5910:

11/3 is correct though As a decimal, it would be roughly 11/3 = 3.67

jimthompson5910:

Function 1 has a slope of 1 Function 2 has a slope of 11/3 = 3.67 Now you select which is larger You'll also need to check the y intercepts of each function as well.

Jaydybird:

function 2 is larger

jimthompson5910:

yes function 2 has the larger slope Now you'll have to compare the y intercepts

Jaydybird:

and how do i do that

jimthompson5910:

In the table, look where x = 0. What is the paired y value for this x value?

Jaydybird:

1

jimthompson5910:

yes, so the y intercept is 1 here For function 2, it's not as quick, but it's not too bad We found the slope to be m = 11/3 We'll use this along with the point (x,y) = (-1,-4) to find the y intercept value b Recall that y = mx+b has m as the slope and b as the y intercept So, y = mx+b -4 = (11/3)*(-1) + b -4 = -11/3 + b 3*(-4) = 3(-11/3 + b) -12 = -11 + 3b I'll let you finish up and solve for b. I multiplied both sides by 3 to clear out the fraction.

Jaydybird:

so if im doing this right 26=b

jimthompson5910:

incorrect

Jaydybird:

sorry -26=b

jimthompson5910:

think of -12 = -11 + 3b as -12 = 3b-11 which is the same as 3b-11 = -12

jimthompson5910:

how would you solve 3b-11 = -12 ?

Jaydybird:

would it be 15?

jimthompson5910:

first we add 11 to both sides, agreed?

Jaydybird:

ya

Jaydybird:

that would make it =-1 right?

jimthompson5910:

That means 3b-11 = -12 becomes 3b-11+11 = -12+11 The -11+11 turns into 0 the -12+11 turns into -1

jimthompson5910:

yes you beat me to it

jimthompson5910:

so we now have 3b = -1

jimthompson5910:

Now we need to fully isolate b

Jaydybird:

then you divide?

jimthompson5910:

yes

Jaydybird:

so would it be -1/3

jimthompson5910:

yep b = -1/3 that's the y intercept of function 2

jimthompson5910:

summary: y intercept of function 1 is 1 y intercept of function 2 is -1/3

Jaydybird:

so function 1 is greater

jimthompson5910:

yes

Jaydybird:

omg thanks

Jaydybird:

you helped me a lot

jimthompson5910:

I'm glad I could help